# Nested ensemble Kalman filter for static parameter inference in nonlinear state-space models

### v1

@@ Abstract
The ensemble Kalman filter (EnKF) is a popular technique for performing inference in state-space models (SSMs), particularly when the dynamic process is high-dimensional. Unlike reweighting methods such as sequential Monte Carlo (SMC, i.e. particle filters), the EnKF leverages either the linear Gaussian structure of the SSM or an approximation thereof, to maintain diversity of the sampled latent states (the so-called ensemble members) via shifting-based updates. Joint parameter and state inference using an EnKF is typically achieved by augmenting the state vector with the static parameter. In this case, it is assumed that both parameters and states follow a linear Gaussian state-space model, which may be unreasonable in practice. In this paper, we combine the reweighting and shifting methods by replacing the particle filter used in the SMC2 algorithm of Chopin et al., with the ensemble Kalman filter. Hence, parameter particles are weighted according to the estimated observed-data likelihood from the latest observation computed by the EnKF, and particle diversity is maintained via a resample-move step that targets the marginal parameter posterior under the EnKF. Extensions to the resulting algorithm are proposed, such as the use of a delayed acceptance kernel in the rejuvenation step and incorporation of nonlinear observation models. We illustrate the resulting methodology via several applications.
@@ 1 Introduction
State-space models (SSMs, e.g. harvey1990forecasting; shumway2006; Sarkka2013) provide a flexible framework for modelling time-series data. An SSM typically involves two components: a model describing the temporal evolution of the latent system state and a model linking the state to observations. When these components involve linear functions of the latent process, the result is a tractable state-space model, often referred to as a dynamic linear model in the discrete-time context (DLM, see e.g. west2006bayesian; petris2009dynamic). In this case, the tractable observed-data likelihood and smoothing process can be efficiently calculated (conditional on parameters) using a Kalman filter (kalman1960) and joint state and parameter inference can proceed in the Bayesian context via Markov chain Monte Carlo (e.g. carter1994gibbs; fruhwirth1994data; GamermanLopes06). For nonlinear state-space models, inference may proceed via pseudo-marginal methods (andrieu+r09) and, in particular, particle MCMC (Andrieu2010).
@@
The focus of this paper is sequential Bayesian inference for the parameters and latent states in nonlinear state-space models. From a state filtering perspective, particle filters (e.g. Gordon1993; Doucet2011; Jacob2015) can be used to recycle posterior samples from one time point to the next through a series of simple reweighting and resampling steps. State and parameter filtering is complicated by the exponential decrease over time in distinct parameter particles. This issue can be mitigated through the use of resample-move (also known as mutation) steps, which rejuvenate parameter particles via a Metropolis-Hastings kernel that has invariant distribution given by the parameter posterior. An example of the resulting scheme in the context of a tractable observed-data likelihood can be found in chopin2002. A pseudo-marginal analogue of this scheme, applicable in the intractable observed-data likelihood context, weights parameter particles by a likelihood estimate obtained from a particle filter over states, and rejuvenates parameter particles via a pseudo-marginal Metropolis-Hastings kernel. The resulting algorithm is termed SMC2 (Chopin2013), due to the nested use of particle filters.
@@
It is well known that particle filters suffer from weight degeneracy for high-dimensional targets (see Agapiou2017, for a discussion). Moreover, the simplest, bootstrap particle filter propagates particles myopically, irrespective of the next observation, which can lead to highly variable weights when observation noise is small relative to the noise in the state process. These problems are addressed, at the expense of some inferential accuracy, by the ensemble Kalman filter (EnKF, evensen1994; see also Katzfuss2016 for an introduction), which has been ubiquitously applied in a field of research originating in the geosciences termed data assimilation (e.g. evensen2022). As with the bootstrap particle filter, the EnKF propagates a set of simulations (known as ensemble members) via forward simulation of the state process. Unlike the reweighting of the particle filter, samples are shifted by a vector whose distribution depends on the next observation by leveraging a Gaussian approximation of the state distribution at the observation time, and if need be, of the observation model.
@@
Compared to the state filtering problem, relatively few works have addressed both state and parameter filtering. Moreover, commonly used approaches are often based on strong assumptions. For example, state augmentation (e.g. anderson01; Evensen2007), assumes a linear Gaussian relationship between parameters, observations and states, with the latter augmented to include all parameter components. Several sequential and batch inference schemes are given in Katzfuss2019 (see also Stroud2018) in the context of dynamic parameters. These are applicable in the static parameter context through the assumption of an artificial evolution kernel, as has also been used in the particle filter context (e.g. liu2001). However, this approach can be sensitive to the choice of tuning parameters (vieira2016).
@@
Our contribution is a scheme for state and parameter filtering in which parameter particles are weighted by an approximate observed-data likelihood contribution, computed using the EnKF. Parameter rejuvenation mutates particles through a Metropolis-Hastings kernel that again uses the EnKF to calculate the likelihood for new parameter values. This corresponds to one iteration of the (offline) ensemble MCMC strategy of Katzfuss2019 and drovandi22. The resulting nested ensemble Kalman filter (NEnKF) uses the same nesting approach as SMC2, albeit with the particle filter over states replaced by the EnKF. Given the robustness of the EnKF over a particle filter in high-dimensional settings, we anticipate that the NEnKF will be of most benefit to practitioners working with models in which the number of static parameter components is small relative to the dimension of the hidden state process. Several examples of such a scenario are given in Katzfuss2019.
As with SMC2, the resample-move step in the NEnKF is the main computational bottleneck, and we therefore aim to improve its efficiency through the use of a delayed-acceptance kernel (e.g. christen2005; Goli15; Banterle19) that first tries proposals using a simple $k$-nearest neighbour approximation to the observed data likelihood, that can be constructed dynamically using the current particle set. We also consider the case of nonlinear observation models, by using the EnKF to propagate the state process inside a particle filter. Unlike related approaches which use the EnKF as a proposal mechanism (e.g. bi15; tamboli2021), our novel method allows for marginalising the latent state process between observation times.
@@ 2.2 Particle filter
The bootstrap particle filter (see e.g. Gordon1993) consists of a sequence of weighted resampling steps, whereby upon receipt of $y_{t}$, $N$ state particles $x_{t-1}^{1:N}=\{x_{t-1}^{1},\ldots,x_{t-1}^{N}\}$ are propagated forward via some proposal density, appropriately weighted and resampled with replacement.
@@
The particle filter can be used to unbiasedly estimate the observed-data likelihood via the product (over time) of the average unnormalised weights (see e.g. DelMoral2004; pitt12). The estimator is
@@
which we let depend explicitly on the particles $x_{0:T}^{1:N}$ and their genealogy $a_{0:T-1}^{1:N}$. In the context of batch inference for the static parameters $\theta$, (2.2) can be used inside a pseudo-marginal Metropolis-Hastings (PMMH) scheme which exactly targets $\pi_{T}(\theta|y_{0:T})$ (see e.g. Andrieu2010). The contributions in (2.2) can also be used as incremental weights in a particle filter over static parameters, which we briefly review in the next section.
@@ 2.3 SMC2
which is complicated by the observed-data likelihood $p_{t}^{(\theta)}(y_{t}|y_{0:t-1})$, which is typically unavailable in closed form. Given a weighted sample of size $M$, $\{\theta^{1:M},u_{t-1}^{1:M}\}$, from $\pi_{t-1}(\theta|y_{0:t-1})$, an idealised SMC scheme generates a weighted sample distributed according to (5) by reweighting each $\theta^{i}$ according to $u_{t}^{i}\propto u_{t-1}^{i}p_{t}^{(\theta^{i})}(y_{t}|y_{0:t-1})$. Triggered by the fulfilment of some degeneracy criterion, such as that related to (6), below, a resample-move step (see e.g. gilks2001) generates new parameter particles via a Metropolis-Hastings kernel that has (5) as its invariant distribution. Full details of the resulting iterated batch importance sampling (IBIS) scheme are given in chopin2002.
We consider here the pseudo-marginal analogue of IBIS (termed SMC2, Chopin2013), wherein each $\theta^{i}$ particle is reweighted according to $u_{t}^{i}\propto u_{t-1}^{i}\widehat{p}_{t}^{(\theta^{i})}(y_{t}|y_{0:t-1},x_{t}^{i,1:N},a_{t-1}^{i,1:N})$, and $\widehat{p}_{t}^{(\theta^{i})}(y_{t}|y_{0:t-1},x_{t}^{i,1:N},a_{t-1}^{i,1:N})$ is the estimate of observed-data likelihood obtained by running Algorithm 1 with parameter $\theta^{i}$ and the associated state particle system $\{x_{t-1}^{i,1:N},w_{t-1}^{i,1:N}\}$. Hence, over the entire observation period, each particle $\theta^{i}$ has associated state particles and genealogies $\{x_{0:T}^{i,1:N},a_{0:T-1}^{i,1:N}\}$.
@@
A formal justification of SMC2 can be found in Chopin2013; briefly, if $\psi_{t}^{(\theta)}(x_{0:t}^{1:N},a_{0:t-1}^{1:N})$ denotes the joint probability density of all random variables generated by the state particle filter up to time $t$, then SMC2 recursively targets
@@
   2. Resample. Normalise the weights to give $u_{t}^{i}=\tilde{u}_{t}^{i}/\sum_{k=1}^{M}\tilde{u}_{t}^{k}$. Sample an index $b_{i}$ from $\{1,\ldots,M\}$ with probabilities $u_{t}^{1:M}$. Put $\{\theta^{i},u_{t}^{i}\}:=\{\theta_{t}^{b_{i}},1/M\}$ and $\{x_{t}^{i,1:N},w_{t}^{i,1:N}\}:=\{x_{t}^{b_{i},1:N},w_{t}^{b_{i},1:N}\}$.
@@
Note that the overall acceptance rate of the mutation step will depend, inter alia, on the variance of the likelihood approximation, which can be controlled by dynamically choosing the number of state particles $N$, for example, by doubling $N$ if the empirical acceptance rate falls below some threshold chosen by the practitioner. For a given $\theta$ particle, state particles and their genealogies $\{x_{0:t}^{1:N},a_{0:t-1}^{1:N}\}$ are replaced by $\{\check{x}_{0:t}^{1:\check{N}},\check{a}_{0:t-1}^{1:\check{N}}\}$ (generated through recursive application of Algorithm 1) using the generalised importance sampling strategy of delMoral06. Implementation of this step implicitly weights each $\theta^{i}$ particle by
@@
where $L_{t}^{(\theta)}$ is an artificial backward kernel. As discussed in Chopin2013, although the optimal backward kernel is typically intractable, $L_{t}^{(\theta)}$ can be chosen to give a simple incremental weight of the form $\widehat{p}_{t}^{(\theta^{i})}(y_{0:t}|\check{x}_{0:t}^{i,1:\check{N}},\check{a}_{0:t-1}^{i,1:\check{N}})/\widehat{p}_{t}^{(\theta^{i})}(y_{0:t}|x_{0:t}^{i,1:N},a_{0:t-1}^{i,1:N})$. On the other hand, botha23 suggest a choice of backward kernel that introduces no additional variability in the particle weights; the incremental weight is 1 in this case.
@@ 3 Inference using the ensemble Kalman filter
In this section, we briefly review several existing methods for sequential state and parameter inference based on the ensemble Kalman filter (EnKF). For a detailed discussion, we refer the reader to Katzfuss2019.
@@ 3.1 Ensemble Kalman Filter
The ensemble Kalman filter (EnKF, see e.g. evensen1994; Katzfuss2016) can be used to give approximate draws from the filtering densities $p_{t}^{(\theta)}(x_{t}|y_{0:t})$, $t=0,1,\ldots,T$. We assume here a linear and Gaussian observation model; extensions of the EnKF to other observation models are considered in Section 4.3.
@@
Although (11) can be sampled from trivially, it is common to perform a stochastic update step (see e.g. Katzfuss2016), which moves ensemble members according to a Gaussian pseudo-observation $\tilde{y}_{t}^{j}\sim\mathcal{N}(H\tilde{x}_{t}^{j},R)$. New ensemble members are obtained as $x_{t}^{j}=\tilde{x}_{t}^{j}+\hat{K}_{t}(y_{t}-\tilde{y}_{t}^{j})$ which avoids draws of dimension $d_{x}$ (with $d_{x}>>d_{y}$ in common applications). Note that marginalising over $x_{t}$ in (8) with respect to (9) gives the denominator in (10) (that is, observed data likelihood contribution) as

### v2

@@ Abstract
The ensemble Kalman filter (EnKF) is a popular technique for performing inference in state-space models (SSMs), particularly when the dynamic process is high-dimensional. Unlike reweighting methods such as sequential Monte Carlo (SMC, i.e. particle filters), the EnKF leverages either the linear Gaussian structure of the SSM or an approximation thereof, to maintain diversity of the sampled latent states (the so-called ensemble members) via shifting-based updates. Joint parameter and state inference using an EnKF is typically achieved by augmenting the state vector with the static parameter. In this case, it is assumed that both parameters and states follow a linear Gaussian state-space model, which may be unreasonable in practice. In this paper, we combine the reweighting and shifting methods by replacing the particle filter used in the SMC2 algorithm of Chopin et al. (2013), with the ensemble Kalman filter. Hence, parameter particles are weighted according to the estimated observed-data likelihood from the latest observation computed by the EnKF, and particle diversity is maintained via a resample-move step that targets the marginal parameter posterior under the EnKF. Extensions to the resulting algorithm are proposed, such as the use of a delayed acceptance kernel in the rejuvenation step and incorporation of nonlinear observation models. We illustrate the resulting methodology via several applications.
@@ 1 Introduction
State-space models (SSMs, e.g. Harvey, 1990; Shumway and Stoffer, 2006; Särkkä, 2013) provide a flexible framework for modelling time-series data. An SSM typically involves two components: a model describing the temporal evolution of the latent system state and a model linking the state to observations. When these components involve linear functions of the latent process, the result is a tractable state-space model, often referred to as a dynamic linear model in the discrete-time context (DLM, see e.g. West and Harrison, 2006; Petris et al., 2009). In this case, the tractable observed-data likelihood and smoothing process can be efficiently calculated (conditional on parameters) using a Kalman filter (Kalman, 1960) and joint state and parameter inference can proceed in the Bayesian context via Markov chain Monte Carlo (e.g. Carter and Kohn, 1994; Frühwirth-Schnatter, 1994; Gamerman and Lopes, 2006). For nonlinear state-space models, inference may proceed via pseudo-marginal methods (Andrieu and Roberts, 2009) and, in particular, particle MCMC (Andrieu et al., 2010).
@@
The focus of this paper is sequential Bayesian inference for the parameters and latent states in nonlinear state-space models. From a state filtering perspective, particle filters (e.g. Gordon et al., 1993; Doucet and Johansen, 2011; Jacob, 2015) can be used to recycle posterior samples from one time point to the next through a series of simple reweighting and resampling steps. State and parameter filtering is complicated by the exponential decrease over time in distinct parameter particles. This issue can be mitigated through the use of resample-move (also known as mutation) steps, which rejuvenate parameter particles via a Metropolis-Hastings kernel that has invariant distribution given by the parameter posterior. An example of the resulting scheme in the context of a tractable observed-data likelihood can be found in Chopin (2002). A pseudo-marginal analogue of this scheme, applicable in the intractable observed-data likelihood context, weights parameter particles by a likelihood estimate obtained from a particle filter over states, and rejuvenates parameter particles via a pseudo-marginal Metropolis-Hastings kernel. The resulting algorithm is termed SMC2 (Chopin et al., 2013), due to the nested use of particle filters.
@@
It is well known that particle filters suffer from weight degeneracy for high-dimensional targets (see Agapiou et al., 2017, for a discussion). Moreover, the simplest, bootstrap particle filter propagates particles myopically, irrespective of the next observation, which can lead to highly variable weights when observation noise is small relative to the noise in the state process. These problems are addressed, at the expense of some inferential accuracy, by the ensemble Kalman filter (EnKF, Evensen, 1994; see also Katzfuss et al., 2016 for an introduction), which has been ubiquitously applied in a field of research originating in the geosciences termed data assimilation (e.g. Evensen et al., 2022). As with the bootstrap particle filter, the EnKF propagates a set of simulations (known as ensemble members) via forward simulation of the state process. Unlike the approach taken by the particle filter, forward samples are shifted, rather than reweighted and subsequently resampled, by leveraging a Gaussian approximation of the state distribution at the observation time, and if need be, of the observation model. Thus, weight collapse is avoided, potentially at the expense of inferential accuracy.
@@
Compared to the state filtering problem, relatively few works have addressed both state and parameter filtering. Moreover, commonly used approaches are often based on strong assumptions. For example, state augmentation (e.g. Anderson, 2001; Evensen, 2009), assumes a linear Gaussian relationship between parameters, observations and states, with the latter augmented to include all parameter components. Several sequential and batch inference schemes are given in Katzfuss et al. (2019) (see also Stroud et al., 2018) in the context of dynamic parameters. These are applicable in the static parameter context through the assumption of an artificial evolution kernel, as has also been used in the particle filter context (e.g. Liu and West, 2001). However, this approach can be sensitive to the choice of tuning parameters (Vieira and Wilkinson, 2016).
@@
Our contribution is a scheme for state and parameter filtering in which parameter particles are weighted by an approximate observed-data likelihood contribution, computed using the EnKF. Parameter rejuvenation mutates particles through a Metropolis-Hastings kernel that again uses the EnKF to calculate the likelihood for new parameter values. This corresponds to one iteration of the (offline) ensemble MCMC strategy of Katzfuss et al. (2019) and Drovandi et al. (2022). The resulting nested ensemble Kalman filter (NEnKF) uses the same nesting approach as SMC2, albeit with the particle filter over states replaced by the EnKF. Given the robustness of the EnKF over a particle filter in high-dimensional settings, we anticipate that the NEnKF will be of most benefit to practitioners working with models in which the number of static parameter components is small relative to the dimension of the hidden state process. Several examples of such a scenario are given in Katzfuss et al. (2019).
As with SMC2, the resample-move step in the NEnKF is the main computational bottleneck. This can be alleviated at the expensive of inferential accuracy, for example, by calculating the likelihood over a moving window rather than the full time horizon (e.g. Del Moral et al., 2017) or by replacing the resample-move step with parameter particles sampled from a jittering kernel (e.g. Crisan and Míguez, 2018; Fang et al., 2024). In this work, we choose to improve the efficiency of the resample-move through the use of a delayed-acceptance kernel (e.g. Christen and Fox, 2005; Golightly et al., 2015; Banterle et al., 2019) that first tries proposals using a simple $k$-nearest neighbour approximation to the observed data likelihood, that can be constructed dynamically using the current particle set.
We also consider the case of nonlinear and non-Gaussian observation models, for which using the EnKF directly can lead to biased or inaccurate estimates of the filtering distribution (e.g. Zhang et al., 2010; Sætrom and Omre, 2011; Grooms, 2022). In this case we propose to use the EnKF to propagate the state process inside the inner particle filter of an SMC2 scheme, so that the bias in the enKF is corrected for via an appropriate weight function. Unlike related approaches which use the EnKF as a proposal mechanism inside a particle filter (e.g. Bi et al., 2015; Tamboli, 2021), our novel method allows for marginalising the latent state process between observation times.
Finally, we note that replacing the inner particle filter of SMC2 with the ensemble Kalman filter has been proposed independently of our work by Temfack and Wyse (2025). Unlike our approach, the main focus of their work is sequential inference for stochastic epidemic models, with the EnKF adapted to epidemic data by incorporating state-dependent observation variance, as is necessary for over-dispersed incidence counts. Additionally, and as in Drovandi et al. (2022), an unbiased Gaussian density estimator is used to mitigate finite-sample effects. The nested use of the EnKF in Temfack and Wyse (2025) further underlines the possible computational gains of using the EnKF versus a vanilla SMC2 implementation.
@@ 2.2 Particle filter
The bootstrap particle filter (see e.g. Gordon et al., 1993) consists of a sequence of weighted resampling steps, whereby upon receipt of $y_{t}$, $N$ state particles $x_{t-1}^{1:N}=\{x_{t-1}^{1},\ldots,x_{t-1}^{N}\}$ are propagated forward via some proposal density, appropriately weighted and resampled with replacement.
@@
The particle filter can be used to unbiasedly estimate the observed-data likelihood via the product (over time) of the average unnormalised weights (see e.g. Del Moral, 2004; Pitt et al., 2012). The estimator is
@@
which we let depend explicitly on the particles $x_{0:T}^{1:N}$ and their genealogy $a_{0:T-1}^{1:N}$. In the context of batch inference for the static parameters $\theta$, (2.2) can be used inside a pseudo-marginal Metropolis-Hastings (PMMH) scheme which exactly targets $\pi_{T}(\theta|y_{0:T})$ (see e.g. Andrieu et al., 2010). The contributions in (2.2) can also be used as incremental weights in a particle filter over static parameters, which we briefly review in the next section.
@@ 2.3 SMC2
which is complicated by the observed-data likelihood $p_{t}^{(\theta)}(y_{t}|y_{0:t-1})$, which is typically unavailable in closed form. Given a weighted sample of size $M$, $\{\theta^{1:M},u_{t-1}^{1:M}\}$, from $\pi_{t-1}(\theta|y_{0:t-1})$, an idealised SMC scheme generates a weighted sample distributed according to (5) by reweighting each $\theta^{i}$ according to $u_{t}^{i}\propto u_{t-1}^{i}p_{t}^{(\theta^{i})}(y_{t}|y_{0:t-1})$. Triggered by the fulfilment of some degeneracy criterion, such as that related to (6), below, a resample-move step (see e.g. Gilks and Berzuini, 2001) generates new parameter particles via a Metropolis-Hastings kernel that has (5) as its invariant distribution. Full details of the resulting iterated batch importance sampling (IBIS) scheme are given in Chopin (2002).
We consider here the pseudo-marginal analogue of IBIS (termed SMC2, Chopin et al., 2013), wherein each $\theta^{i}$ particle is reweighted according to $u_{t}^{i}\propto u_{t-1}^{i}\widehat{p}_{t}^{(\theta^{i})}(y_{t}|y_{0:t-1},x_{t}^{i,1:N},a_{t-1}^{i,1:N})$, and $\widehat{p}_{t}^{(\theta^{i})}(y_{t}|y_{0:t-1},x_{t}^{i,1:N},a_{t-1}^{i,1:N})$ is the estimate of observed-data likelihood obtained by running Algorithm 1 with parameter $\theta^{i}$ and the associated state particle system $\{x_{t-1}^{i,1:N},w_{t-1}^{i,1:N}\}$. Hence, over the entire observation period, each particle $\theta^{i}$ has associated state particles and genealogies $\{x_{0:T}^{i,1:N},a_{0:T-1}^{i,1:N}\}$.
@@
A formal justification of SMC2 can be found in Chopin et al. (2013); briefly, if $\psi_{t}^{(\theta)}(x_{0:t}^{1:N},a_{0:t-1}^{1:N})$ denotes the joint probability density of all random variables generated by the state particle filter up to time $t$, then SMC2 recursively targets
@@
   2. Resample. Normalise the weights to give $u_{t}^{i}=\tilde{u}_{t}^{i}/\sum_{k=1}^{M}\tilde{u}_{t}^{k}$. Sample an index $b_{i}$ from $\{1,\ldots,M\}$ with probabilities $u_{t}^{1:M}$. Put $\{\theta^{i},u_{t}^{i}\}:=\{\theta^{b_{i}},1/M\}$ and $\{x_{t}^{i,1:N},w_{t}^{i,1:N}\}:=\{x_{t}^{b_{i},1:N},w_{t}^{b_{i},1:N}\}$.
@@
Note that the overall acceptance rate of the mutation step will depend, inter alia, on the variance of the likelihood approximation, which can be controlled by dynamically choosing the number of state particles $N$, for example, by doubling $N$ if the empirical acceptance rate falls below some threshold chosen by the practitioner. For a given $\theta$ particle, state particles and their genealogies $\{x_{0:t}^{1:N},a_{0:t-1}^{1:N}\}$ are replaced by $\{\check{x}_{0:t}^{1:\check{N}},\check{a}_{0:t-1}^{1:\check{N}}\}$ (generated through recursive application of Algorithm 1) using the generalised importance sampling strategy of Del Moral et al. (2006). Implementation of this step implicitly weights each $\theta^{i}$ particle by
@@
where $L_{t}^{(\theta)}$ is an artificial backward kernel. As discussed in Chopin et al. (2013), although the optimal backward kernel is typically intractable, $L_{t}^{(\theta)}$ can be chosen to give a simple incremental weight of the form $\widehat{p}_{t}^{(\theta^{i})}(y_{0:t}|\check{x}_{0:t}^{i,1:\check{N}},\check{a}_{0:t-1}^{i,1:\check{N}})/\widehat{p}_{t}^{(\theta^{i})}(y_{0:t}|x_{0:t}^{i,1:N},a_{0:t-1}^{i,1:N})$. On the other hand, Botha et al. (2023b) suggest a choice of backward kernel that introduces no additional variability in the particle weights; the incremental weight is 1 in this case.
@@ 3 Inference using the ensemble Kalman filter
In this section, we briefly review several existing methods for sequential state and parameter inference based on the ensemble Kalman filter (EnKF). For a detailed discussion, we refer the reader to Katzfuss et al. (2019).
@@ 3.1 Ensemble Kalman Filter
The ensemble Kalman filter (EnKF, see e.g. Evensen, 1994; Katzfuss et al., 2016) can be used to give approximate draws from the filtering densities $p_{t}^{(\theta)}(x_{t}|y_{0:t})$, $t=0,1,\ldots,T$. We assume here a linear and Gaussian observation model; extensions of the EnKF to other observation models are considered in Section 4.3.
@@
Although (11) can be sampled from trivially, it is common to perform a stochastic update step (see e.g. Katzfuss et al., 2016), which moves ensemble members according to a Gaussian pseudo-observation $\tilde{y}_{t}^{j}\sim\mathcal{N}(H\tilde{x}_{t}^{j},R)$. New ensemble members are obtained as $x_{t}^{j}=\tilde{x}_{t}^{j}+\hat{K}_{t}(y_{t}-\tilde{y}_{t}^{j})$ which avoids draws of dimension $d_{x}$ (with $d_{x}>>d_{y}$ in common applications). Note that marginalising over $x_{t}$ in (8) with respect to (9) gives the denominator in (10) (that is, observed data likelihood contribution) as